Reports 1-1 of 1 Clear search Modify search
AOS (Baffles & Beam dumps)
tomotada.akutsu - 21:48 Thursday 28 July 2022 (21638) Print this report
Invac work finalization in IFI and IMM chambers: Day 3

N. Satō, Akutsu; following 21578


Figs. 1 and 2 are updates from the past summary in 21564 and 21578; completed cable management for thermometers relavant to the high-power beam dumps; continued cares for stray light.

Thermometer cabling for invac high-power beam dump

As touched in 21009 and 20963, there have been homework regarding the high-power beam dump system, and we completed it.

  • Today we set an additional thermometer to monitor the temperatre of the top surface of the optical bench in the IFI chamber (Figs. 3 and 4).
  • We also introduced Dsub anchors to hold cables stably in the chamber (Fig. 5).

Currently all the three thermometers are read by a cryocon located in the 2nd floor of the clean booth above the IFI chamber. Additional some work would be needed to read the thermometer data from the epics channels.

Beam shield for IMMT1 coming back

Three beam shield plates around the IMMT1 suspension structure were detached (16920) last year. So we got them back with black-head screws and washers (Fig. 6). We also collected two leftover positioners (Fig. 7), which are incompatible with vacuum.

Replacing with SiC for beam dump #6 in IMM

Beam dump #6 in the IMM chamber has been once broken during O3GK maybe due to accidental hitting of high power beam, and so detached (16920), and then got back with still KG5 plates (19588), but anticipating these KG5 glasses will be replaced with SiC plates. We are not so much sure if this would be the main cause of the break of KG5 of this beam dump, though

Anyway, today we replaced these KG5 glasses with single-side-polished SiC plates (Fig. 8; the mechanical dimension is JGW-D1504158, while the one-side polish was Ra0.01 according to the ordering e-mails). The polished surfaces are used for the inner sides of the V-shape beam dump to trap the incident ghost beam with aniticipating small back-scattering of incident light. Hopefully this time, thanks to the high heat conductivity of SiC, such break would not happen.


  • Usually, looking at the KG5 surface with an IR viewer, no obvious IR spots could not be seen, but in Fig. 8 there are. This may be partly due to that the incident light beam power might be too large, but maybe partly due to the surface-ish scattering may be larger for (even though polished) SiC than that of raw KG5 plate, maybe (?); I am not well sure.
  • With a small piece of a sensor card, I could check whether the 1st ghost beam of forward beam at IMMT1 would be caught by the beam dump behind IMMT1, but sadly, the answer was negative. This should not be allowed left, so we will do some countermeasure tomorrow.
Images attached to this report
Comments to this report:
koji.nakagaki - 15:10 Wednesday 10 August 2022 (21750) Print this report

The IP address and other settings were configured on the IFI CRYOCON and connected to the network.


koji.nakagaki - 17:48 Friday 19 August 2022 (21835) Print this report
[Ikeda, Nakagaki]

EPICS Channels for IFI Cryocon was set up and the update process was registered in k1script1.

The following was set
takaaki.yokozawa - 11:30 Monday 22 August 2022 (21843) Print this report
The 3day trend.
(There are small gap in CHA and CHB???)
Images attached to this comment
Search Help